Suppose the temperature of a body when discovered is 85o F. Two hours later, the 

temperature is 74o F and the room temperature is 68o F. Find the time when the body was discovered after death (assume the body temperature to be 98.6o F at the time of death.)


1
Expert's answer
2021-06-03T12:39:06-0400

Newton's law of cooling states that the temperature of a body changes at a rate proportional to the difference in temperature between the body and its surroundings.

Let "T_0" be the initial temperature of a body and "T" be the temperature of the body at time "t" . If "T_s" be the temperature of the surroundings, then Newton's law of cooling states that

"\frac{dT}{dt}=-k(T-T_s)" ,

where "k" is a positive constant.

"\Rightarrow \frac{dT}{ T-T_s}=-kdt"

Integrating both sides,

"\int_{T_0}^T \frac{dT}{ T-T_s}=-k\int _0^tdt"

"\Rightarrow[ ln(T-T_s)]_{T_0}^T=-kt\\
\Rightarrow ln(\frac{T-T_s}{T_0-T_s})=-kt\\
\Rightarrow \frac{T-T_s}{T_0-T_s}=e^{-kt}\ ...............(1)"

In the given problem, initial temperature "T_0=98.6\degree F" ( temperature of the body at the time of death "t=0)"

and surrounding temperature "T_s=68\degree F" ( room temperature).

Let at "t=t_1" the body was discovered and its temperature was "T_1=85\degree F"

From Eq.(1),

"\frac{T_1-T_s}{T_0-T_s}=e^{-kt_1}\ ...............(2)"

After two hours at "t=t_1+2" the temperature of the body is "T_2=74\degree F"

From Eq.(1),

"\frac{T_2-T_s}{T_0-T_s}=e^{-k(t_1+2)}\ ...............(3)"

Dividing Eq.(2) by Eq.(3),

"\frac{T_1-T_s}{T_2-T_s}=\frac{e^{-kt_1}}{e^{-k(t_1+2)}}\\
\Rightarrow \frac{T_1-T_s}{T_2-T_s}=e^{2k}\\
\Rightarrow k=\frac{1}{2}ln(\frac{T_1-T_s}{T_2-T_s})\\
\Rightarrow k=\frac{1}{2}ln(\frac{85-68}{74-68})\\
\Rightarrow k=0.52"

Substitute the values of "k, T_1, T_s, T_0" in Eq.(2), we will get "t_1" .

"\therefore \frac{85-68}{98.6-68}=e^{-0.52t_1}"

On simplification, "t_1=1.13\ hours"

Therefore, the body was discovered 1.13 hours after death.